OpenStudy (anonymous):

Simplify and select the answer with the appropriate restrictions for the variables. 10x^2-15x/5x^2 A.2-15x,(x≠0) B.2-3x,(x≠0) C.2x-3/x,(x≠0) D.10x-3,(x≠0)

OpenStudy (anonymous):

10x^3 -3

OpenStudy (anonymous):

i dont think thats right but thank you! :)

OpenStudy (anonymous):

it's math not think you simplify 15/5=3 x/x^2=1/x multiple by x you find it

OpenStudy (anonymous):

ohh okay thank you i was like...okay yea thank you

OpenStudy (anonymous):

i think its c but idk for sure

OpenStudy (anonymous):

thank you!